(7) Reduce the following complex fractions: 1/5 0.1 (8) GROWTH RATES (RATES OF CHANGE Country 100 million 110 milion 1999 population 2,000 million 2,040 million 2000 population Amount of change in population Percentage change in population mind = amount0fchange in population x 100 amount of original population (a) Which country had the largest increase in its population? (b) Which country had the largest percentage increase in its population? So even though population grew by a much larger amount in country # -it has a lower growth rate (or percentage rate of change). The population is increasing in both, but it is increasing more slowly in one country
Answer #1

Solution:

Country 1

Country 2

1999 population

100 million

2,000 million

2000 population

110 million

2,040 million

Amount of change in population

10 million

40 million

Working

110 - 100 = 10

2040 - 2000 = 40

Percentage change in population

10%

2%

Working

10/100*100 = 10%

40/2000*100 = 2%

1) Country 2 has the largest increase in population. Calculated above.

2) Country 2 has the largest percentage increase in population. Calculated above.

3) Country 2

Because the population grew higher in Country 2. Calculated above.